# Learning to Learn with Contrastive Meta-Objective

### v3

@@ 1 Introduction
Learning to learn, also known as meta-learning (schmidhuber1987evolutionary, ; thrun1998learning, ), is a powerful paradigm designed to enable learning systems to adapt quickly to new tasks. During the meta-training phase, a meta-learner simulates adaptation (learning) across a variety of relevant tasks to accumulate knowledge on how to learn effectively. In the meta-testing phase, this learned adaptation strategy is applied to unseen tasks. The adaptation is typically accomplished by the meta-learner, which, given a set of task-specific training examples, generates a predictive model tailored to that task. Meta-learning has been widely applied to important learning problems such as few-shot learning (finn2017model, ; wang2020generalizing, ), reinforcement learning (yu2020meta, ; nagabandi2019learning, ), and neural architecture search (elsken2020meta, ; shaw2019meta, ).
@@
The mini-batch episodic training framework is universal, and naturally gives the information about task identity, which can serve as additional supervision for meta-training for generalizability. Inspired by the intrinsic property of human’s fast learning ability: alignment and discrimination (chen2012object, ; hummel2013object, ; christie2021learning, ), we hope a meta-learner itself should be able to tell if different datasets are from the same task or different tasks by exploiting task identity. A good learner possesses the alignment ability to align different partial views of a certain object, which means they can integrate various aspects or perspectives of information to form a coherent understanding (christian2021alignment, ). This means a meta-learner should learn similar models from different datasets of the same tasks even if the data are few or noisy, benefiting the meta-testing performance through being robust to the given labeled data. Discrimination involves distinguishing between similar stimuli to respond appropriately only to decisive inputs (robbins1970stimulus, ). This means a meta-learner should learn different models from different tasks even if some of their inputs are similar, benefiting meta-testing performance through generalization to diverse tasks.
@@
ConML distinguishes itself by being universal: it is problem-agnostic, as it is based-on mini-batch episodic training where task-identity are intrinsic information; and it is learner-agnostic, as we design easy-to-implement mapping functions from model to representations for different meta-learners. Additionally, it is efficient in that it requires no additional data or retraining. Existing approaches have also leveraged task-level alignment or contrastiveness as additional supervision for improved meta-learning. However, they are based on either problem-specific knowledge (wang2017learning, ; ye2022few, ; wang2023robust, ; li2024metacl, ) or learner-specific knowledge (gondal2021function, ; fei2021melr, ). Thus, they can be improved by exploiting the problem- and learner- agnostic task identity through incorporating with ConML (see details in Section 5). Our contributions are:
@@ 2 Preliminaries: Learning to Learn
Different meta-learners implement their own specific functions within $g(;\theta)$. Popular meta-learning approaches can be broadly categorized into the following types (bronskill2021memory, ): (i) Optimization-based approaches (andrychowicz2016learning, ; finn2017model, ; nichol2018first, ), which focus on learning better optimization strategies for adapting to new tasks; (ii) Metric-based approaches (vinyals2016matching, ; snell2017prototypical, ; sung2018learning, ), which leverage learned similarity metrics; and (iii) Amortization-based approaches (garnelo2018conditional, ; requeima2019fast, ; bateni2020improved, ), which aim to learn a shared representation across tasks, amortizing the adaptation process by using neural networks to directly infer task-specific parameters from the training set; (iv) Furthermore, the emerging ICL ability in LLMs can also be viewed as the consequence of meta-learning garg2022can ; akyurek2023learning , and ICL model is meta-learner with minimal inductive bias (wu2025context, ), so we will also use meta-learner $g$ to express the function of ICL. Details reformulating ICL model as meta-learner to incorporate ConML with are provided in Section 3.3.
@@ 3 Meta-Learning with ConML
Now, we introduce our ConML which equips meta-learners with the desired alignment and discrimination ability via task-level contrastive learning. We first present the general framework of ConML based on episodic training in Section 3.1, followed by specifications for the three main streams of meta-learning approaches and ICL in Section 3.3.
@@ 3.1 A General Framework
To enhance the alignment and discrimination abilities of meta-learning, we draw inspiration from Contrastive Learning (CL) (oord2018representation, ; chen2020simple, ; wang2020understanding, ). CL focuses on learning representations that are invariant to irrelevant details while preserving essential information. This is achieved by maximizing alignment and discrimination (uniformity) in the representation space (wang2020understanding, ). While most existing studies focus on sample-wise contrastive learning in the representation space via unsupervised learning (wu2018unsupervised, ; hjelm2018learning, ; bachman2019learning, ; tian2020contrastive, ; he2020momentum, ; chen2020simple, ; khosla2020supervised, ), we extend CL to the model space in meta-learning.
@@
Obtaining Model Representation.  To train the meta-learner $g$, the inner-task distance $d^{\text{in}}$ and inter-task distance $d^{\text{out}}$ are measured in the output space of $g$, also referred to as the model space $\mathcal{H}$. A practical approach is to represent the model $h=g(\mathcal{D};\theta)\in\mathcal{H}$ as a fixed-length vector $\bm{e}\in\mathbb{R}^{d}$, and then compute the distances using an explicit distance function $\phi(\cdot,\cdot)$ (e.g., cosine distance). To generalize across different learners to form a learner-agnostic framework, we introduce a projection function $\psi:\mathcal{H}\rightarrow\mathbb{R}^{d}$ to obtain the model representations $\bm{e}=\psi(h)$, The details of $\mathcal{H}$ and $\psi$ will be specified in Section 3.3.
@@
Obtaining Inter-Task Distance.  Discrimination is achieved by maximizing inter-task distance. Since the goal of meta-learning is to improve performance on unseen tasks, it is crucial for the meta-learner $g$ to generalize well across diverse tasks. Given the natural assumption that different tasks require distinct task-specific models, it is essential that $g$ can learn to differentiate between tasks—i.e., possess strong discrimination capabilities. To enhance task-level generalization, we define the inter-task distance $d^{\text{out}}$, which should be maximized to encourage $g$ to learn distinct models for different tasks. Specifically, for any two tasks $\tau\neq\tau^{\prime}$ during meta-training, we aim to maximize the distance between their respective representations, $\bm{e}^{*}_{\tau}$ and $\bm{e}^{*}_{\tau^{\prime}}$. To make this practical within the mini-batch episodic training paradigm, we compute $d^{\text{out}}$ across a batch of tasks sampled in each episode:
@@
Training Procedure.
ConML optimizes the combination of the original episodic loss $\mathcal{L}_{e}$ and contrastive meta-objective $\mathcal{L}_{c}=d^{\text{in}}-d^{\text{out}}$:
@@
The meta-training procedure with ConML is provided in Algorithm 2. Note ConML is agnostic to the form of $\mathcal{L}_{e}$ so here we take the typical validation loss as example. Compared to Algorithm 1, ConML introduces additional computations for $\psi(g(\mathcal{D};\theta))$ a total of $K+1$ times per episode. However, $\psi$ is implemented as a lightweight function (e.g., extracting model weights), and $g(\mathcal{D};\theta)$ is already part of the standard episodic training process, with multiple evaluations of $g(\mathcal{D};\theta)$ being parallelizable. As a result, ConML incurs only a little extra cost in computation (detailed analysis is in Appendix A).
@@ 3.2 Provable Benefits for Generalization
Here we provide another perspective to understand how ConML helps meta-learning. It is provable that a meta-learner which minimizes $\mathcal{L}_{c}$ has lower generalization error upper-bound than any other meta-learners. This means ConML serves as a ’safeguard’ for the worst case of error due to finite samples in $\mathcal{D}^{\text{val}}_{\tau}$ in meta-testing, that can be plugged-in any meta-learners.
@@
$\!\!\Delta\epsilon_{p(\tau)}(\theta)\!=\!{E_{\tau\sim p(\tau)}}{E_{D_{\tau}^{tr}\sim\tau^{n}}}{E_{(x,y)\sim\tau}}\big[\ell(g(D_{\tau}^{tr};\theta)(x),y)\big]\!-\!\min_{\theta}\!{E_{\tau\sim p(\tau)}}\Big[\min_{h\in H_{g(\theta)}}{\!\!E_{(x,y)\sim\tau}}\big[\ell(h(x),y)\big]\Big],\!\!$
where $H_{g(\theta)}$ is the hypothesis class of $h$ given $g(;\theta)$. The value of $\Delta\epsilon_{{p(\tau)}}(\theta)>0$, means the difference between expectation of validation loss between $g(;\theta)$ given finite $n$ examples per task, and the best we can find given $g$ and $p(\tau)$. First, we can find an upper bound $U_{p(\tau)}(\theta)$ for $\Delta\epsilon_{p(\tau)}(\theta)$ as follows:
@@
Note that typical meta-objective $\mathcal{L}_{e}$ (determined by $g(;\theta)$ and $p(\tau)$) can not provide effective supervision for the value of $U_{p(\tau)}(\theta)$, as the following lemma:
Lemma 2.
$\forall\varepsilon\in\mathbb{R},~\exists p(\tau)$, $\min_{\theta}U_{p(\tau)}(\theta)-U_{p(\tau)}(\theta^{*}_{\mathcal{L}_{e}})\!>\!\varepsilon$, where $\theta^{*}_{\mathcal{L}_{e}}\!=\!\arg\min_{\theta}\mathcal{L}_{e}(g(;\theta),p(\tau))$.
However, given contrastive meta-objective $\mathcal{L}_{c}$ as defined above, we have the following theorem:
@@
This means while typical meta-objective $\mathcal{L}_{e}$ can not guarantee the worst-case meta-testing performance, the contrastive meta-objective can exactly serve as a surrogate objective of the worst case, as described in above theorem. Note that this holds for any $p(\tau)$ and $g$, which indicates the problem- and learner-agnostic benefit of ConML. The proof is provided in Appendix B.
@@ 3.3 Integrating with Typical Meta-Learners
With In-Context Learning (ICL).  An ICL model makes task-specific prediction by $g([\vec{\mathcal{D}},x];\theta)$, where $g$ is a sequential model and $\vec{\mathcal{D}}$ is the sequentialized $\mathcal{D}$ (prompt), $x_{1},y_{1},\cdots,x_{m},y_{m}]$. The detail is provided in Appendix D. Note that ICL does not specify an explicit output model $h(x)=g(\mathcal{D};\theta)(x)$; instead, this procedure exists only implicitly through the feeding-forward of the sequence model. Thus, obtaining the representation $\psi(g(\mathcal{D};\theta))$ by explicit model weights of $h$ is not feasible for ICL. To represent what $g$ learns from $\mathcal{D}$, we design to incorporate $\vec{\mathcal{D}}$ with a dummy input $u$, which functions as a probe and its corresponding output can be readout as representation:
@@
where $u$ is constrained to be in the same shape as $x$, and has consistent value in an episode. For example, for training a ICL model on linear regression tasks we can choose $u=\bm{1}$, and in pre-training of LLM we can choose u=u="what is this task?". The complete algorithm of ConML for training an ICL model is provided in Appendix C.
@@ 4 Empirical Studies
We provide empirical studies in this section. The empirical study to understand the effect of ConML on synthetic data, which shows that learning to learn with ConML brings generalizable alignment and discrimination abilities, and alignment enhances fast-adaptation and discrimination enhances task-level generalizability, is provided in Appendix E. Code is provided at https://github.com/ovo67/ConML_Code.
4.1 Few-Shot Image Classification
We consider representative meta-learning algorithms from different categories, including optimization-based: MAML (finn2017model, ), FOMAML (finn2017model, ), Reptile (nichol2018first, ); metric-based: MatchNet (vinyals2016matching, ), ProtoNet (snell2017prototypical, ); amortization-based: SCNAPs (Simple CNAPS) (bateni2020improved, ); and the state-of-the-art ICL-based few-shot learner: CAML (fifty2024context, ). Note that for CAML, ConML only effect the meta-training of the ICL mode, not the pre-training of Vit feature extractor. We also incorporate ConML with meta-learners with improved meta-training objective as discussed in Section 5, including: MELR (fei2021melr, ) and LastShot (ye2022few, ). We evaluate the meta-learning performance of each algorithm in its original form (w/o ConML) and after incorporating ConML into the training process (w/ ConML). The implementation of ConML follows the general procedure described in Algorithm 2 and the specification for corresponding category in Section 3.3.
@@ 4.1 Few-Shot Image Classification
Table 3: Cross-domain experiment results on META-DATASET (accuracy (%)).
@@
4.2 Cross-Domain Few-Shot Image Classification
To show that ConML is problem-agnostic, we provide learner-agnostic improvement on large-scale cross-domain few-shot image classification problem here. META-DATASET (triantafillou2020meta, ) is a large-scale cross-domain few-shot image classification dataset. Table 3 shows results on META-DATASET. Note that the backbone and setting of P>M>F (hu2022pushing, ) is different with the other baselines (triantafillou2020meta, ), so they are not comparable across baselines. ConML is introduced with the same setting as Section 4.1 (inner-task sampling $K=1$ and $\pi_{\kappa}(\mathcal{D}^{\text{tr}}_{\tau}\cup\mathcal{D}^{\text{val}}_{\tau})=\mathcal{D}^{\text{tr}}_{\tau}$, $\phi(a,b)=1-\nicefrac{{a\cdot b}}{{\|a\|\|b\|}}$ (cosine distance) and $\lambda=0.1$.). Note that for P>M>F, ConML is plugged-in the meta-training phase while keep the other phases unchanged. It can be observed that ConML also brings consistent improvement.
To further show that ConML is problem-agnostic, a new state-of-the-art method on few-shot molecular property prediction problems, provided in Appendix G.

### v4

@@ 1 Introduction
Learning to learn, also known as meta-learning (schmidhuber1987evolutionary, ; thrun1998learning, ), is a powerful paradigm designed to enable learning systems to adapt quickly to new tasks. During the meta-training phase, a meta-learner simulates adaptation (learning) across a variety of relevant tasks to accumulate knowledge on how to learn effectively. In the meta-testing phase, this learned adaptation strategy is applied to unseen tasks. The adaptation is typically accomplished by the meta-learner, which, given a set of task-specific training examples, generates a predictive model tailored to that task.
@@
The mini-batch episodic training framework is universal, and naturally gives the information about task identity, which can serve as additional supervision for meta-training for generalizability. Inspired by the intrinsic property of human’s fast learning ability: alignment and discrimination (chen2012object, ; hummel2013object, ; christie2021learning, ), we hope a meta-learner itself should be able to tell if different datasets are from the same task or different tasks by exploiting task identity. A good learner possesses alignment ability to align different partial views of a certain object, which means they can integrate various aspects or perspectives of information to form a coherent understanding (christian2021alignment, ). This means a meta-learner should learn similar models from different datasets of the same tasks even if the data are few or noisy, benefiting the meta-testing performance through being robust to the given labeled data. Discrimination involves distinguishing between similar stimuli to respond appropriately only to decisive inputs. This means a meta-learner should learn different models from different tasks even if some of their inputs are similar, benefiting meta-testing performance through generalization to diverse tasks.
@@
ConML distinguishes itself by being universal: it is problem-agnostic, as it is based-on mini-batch episodic training where task-identity are intrinsic information; and it is learner-agnostic, as we design easy-to-implement mapping functions from model to representations for different meta-learners. Additionally, it is efficient in that it requires no additional data or retraining. Existing approaches have also leveraged task-level alignment or contrastiveness as additional supervision for improved meta-learning. However, they are based on either problem-specific knowledge (wang2017learning, ; ye2022few, ; wang2023robust, ) or learner-specific knowledge (gondal2021function, ; fei2021melr, ). Thus, they can be improved by exploiting the problem- and learner- agnostic task identity through incorporating with ConML. Our contributions are:
@@ 2 Preliminaries: Learning to Learn
Different meta-learners implement their own specific functions within $g(;\theta)$. Popular meta-learning approaches can be broadly categorized into the following types (bronskill2021memory, ): (i) Optimization-based approaches (andrychowicz2016learning, ; finn2017model, ; nichol2018first, ), which focus on learning better optimization strategies for adapting to new tasks; (ii) Metric-based approaches (vinyals2016matching, ; snell2017prototypical, ; sung2018learning, ), which leverage learned similarity metrics; and (iii) Amortization-based approaches (garnelo2018conditional, ; requeima2019fast, ; bateni2020improved, ), which aim to learn a shared representation across tasks, amortizing the adaptation process by using neural networks to directly infer task-specific parameters from the training set; (iv) Furthermore, the emerging ICL ability in large language models (LLMs) can also be viewed as the consequence of meta-learning garg2022can ; akyurek2023learning , and ICL model is meta-learner with minimal inductive bias (wu2025context, ), so we will also use meta-learner $g$ to express the function of ICL. Details reformulating ICL model as meta-learner to incorporate ConML with are in Section 3.3.
@@ 3 Meta-Learning with ConML
Now, we introduce our ConML which equips meta-learners with the desired alignment and discrimination ability via task-level contrastive learning.
@@ 3.1 A General Framework
To enhance the alignment and discrimination abilities of meta-learning, we draw inspiration from Contrastive Learning (CL) (oord2018representation, ; chen2020simple, ; wang2020understanding, ). CL focuses on learning representations that are invariant to irrelevant details while preserving essential information. This is achieved by maximizing alignment and discrimination (uniformity) in the representation space (wang2020understanding, ). While most existing studies focus on sample-wise contrastive learning in the representation space via unsupervised learning (hjelm2018learning, ; bachman2019learning, ; tian2020contrastive, ; chen2020simple, ), we extend CL to the model space in meta-learning.
@@
Obtaining Model Representation.  To train the meta-learner $g$, the inner-task distance $d^{\text{in}}$ and inter-task distance $d^{\text{out}}$ are measured in the output space of $g$, also referred to as the model space $\mathcal{H}$. A practical approach is to represent the model $h=g(\mathcal{D};\theta)\in\mathcal{H}$ as a fixed-length vector $\bm{e}\in\mathbb{R}^{d}$, and then compute the distances using an explicit distance function $\phi(\cdot,\cdot)$ (e.g., cosine distance). To form a learner-agnostic framework, we introduce a projection function $\psi:\mathcal{H}\rightarrow\mathbb{R}^{d}$ to obtain the model representations $\bm{e}=\psi(h)$. The details of $\mathcal{H}$ and $\psi$ will be specified in Section 3.3.
@@
Obtaining Inter-Task Distance.  Discrimination is achieved by maximizing inter-task distance. Since the goal of meta-learning is to improve performance on unseen tasks, it is crucial for the meta-learner $g$ to generalize well across diverse tasks. Given the natural assumption that different tasks require distinct task-specific models, it is essential that $g$ can learn to differentiate between tasks—i.e., possess strong discrimination capabilities. To enhance task-level generalization, we define the inter-task distance $d^{\text{out}}$, which should be maximized to encourage $g$ to learn distinct models for different tasks. Specifically, for any two tasks $\tau\neq\tau^{\prime}$ during meta-training, we maximize the distance between their respective representations, $\bm{e}^{*}_{\tau}$ and $\bm{e}^{*}_{\tau^{\prime}}$. To make this practical within the mini-batch episodic training paradigm, we compute $d^{\text{out}}$ across a batch of tasks sampled in each episode:
@@
Training Procedure. ConML optimizes the combination of the original episodic loss $\mathcal{L}_{e}$ and contrastive meta-objective $\mathcal{L}_{c}=d^{\text{in}}-d^{\text{out}}$:
@@
The meta-training procedure with ConML is in Algorithm 2. Note ConML is agnostic to the form of $\mathcal{L}_{e}$ so here we take the typical validation loss as example. Compared to Algorithm 1, ConML introduces additional computations for $\psi(g(\mathcal{D};\theta))$ a total of $K+1$ times per episode. However, $\psi$ is implemented as a lightweight function (e.g., extracting model weights), and $g(\mathcal{D};\theta)$ is already part of the standard episodic training process, with multiple evaluations of $g(\mathcal{D};\theta)$ being parallelizable. As a result, ConML incurs only a little extra cost in computation (detailed analysis is in Appendix A).
@@ 3.2 Provable Benefits for Generalization
Here, we provide another perspective to understand how ConML helps meta-learning. It is provable that a meta-learner which minimizes $\mathcal{L}_{c}$ has lower generalization error upper-bound than any other meta-learners. This means ConML serves as a ‘safeguard’ for the worst case of error due to finite samples in $\mathcal{D}^{\text{val}}_{\tau}$ in meta-testing, that can be plugged-in any meta-learners.
@@
$\Delta\epsilon_{p(\tau)}(\theta)\!=\!{E_{\tau\sim p(\tau)}}{E_{D_{\tau}^{tr}\sim\tau^{n}}}{E_{(x,y)\sim\tau}}\ell\big(g(D_{\tau}^{tr};\theta)(x),y\big)\!\!-\!\!\min_{\theta}\!{E_{\tau\sim p(\tau)}}\Big[\!\min_{h\in H_{g(\theta)}}{\!\!\!\!E_{(x,y)\sim\tau}}\ell\big(h(x),y\big)\Big],\!\!$
where $H_{g(\theta)}$ is the hypothesis class of $h$ given $g(;\theta)$. The value of $\Delta\epsilon_{{p(\tau)}}(\theta)>0$, means the difference between expectation of validation loss between $g(;\theta)$ given finite $n$ examples per task, and the best we can find given $g$ and $p(\tau)$. First, we can find an upper bound $U_{p(\tau)}(\theta)$ for $\Delta\epsilon_{p(\tau)}(\theta)$.
@@
Given contrastive meta-objective $\mathcal{L}_{c}$ as defined above, we then have the following theorem:
@@
This means the contrastive meta-objective can exactly serve as a surrogate objective of the worst-case meta-testing performance, as described in the above theorem. Note that this holds for any $p(\tau)$ and $g$, which indicates the problem- and learner-agnostic benefit of ConML. The proof is in Appendix B.
@@ 3.3 Integrating with Typical Meta-Learners
With In-Context Learning (ICL).  An ICL model makes task-specific prediction by $g([\vec{\mathcal{D}},x];\theta)$, where $g$ is a sequential model and $\vec{\mathcal{D}}$ is the sequentialized $\mathcal{D}$ (prompt), $x_{1},y_{1},\cdots,x_{m},y_{m}]$. The details are in Appendix D. Note that ICL does not specify an explicit output model $h(x)=g(\mathcal{D};\theta)(x)$; instead, this procedure exists only implicitly through the feeding-forward of the sequence model. Thus, obtaining the representation $\psi(g(\mathcal{D};\theta))$ by explicit model weights of $h$ is not feasible for ICL. To represent what $g$ learns from $\mathcal{D}$, we design to incorporate $\vec{\mathcal{D}}$ with a dummy input $u$, which functions as a probe and its corresponding output can be readout as representation:
@@
where $u$ is constrained to be in the same shape as $x$, and has consistent value in an episode. For example, for training a ICL model on linear regression tasks we can choose $u=\bm{1}$, and in pretraining of LLM we can choose u=u="what is this task?". The complete algorithm of ConML for training an ICL model is in Appendix C.
@@ 4 Empirical Studies
We provide empirical studies to understand the effect of ConML on synthetic data, which shows that learning to learn with ConML brings generalizable alignment and discrimination abilities. Code is avaliable at https://github.com/LARS-research/ConML.
@@
4.1 Few-Shot Image Classification Performance
We consider representative meta-learning algorithms from different categories, including optimization-based: MAML (finn2017model, ), FOMAML (finn2017model, ), Reptile (nichol2018first, ); metric-based: MatchNet (vinyals2016matching, ), ProtoNet (snell2017prototypical, ); amortization-based: SCNAPs (Simple CNAPS) (bateni2020improved, ); and the state-of-the-art ICL-based few-shot learner: CAML (fifty2024context, ). Note that for CAML, ConML only effect the meta-training of the ICL mode, not the pretraining of Vit feature extractor. We also incorporate ConML with meta-learners with improved meta-training objective as discussed in Section 5, including: MELR (fei2021melr, ) and LastShot (ye2022few, ). We evaluate the meta-learning performance of each algorithm in its original form (w/o ConML) and after incorporating ConML into the training process (w/ ConML). The implementation of ConML follows the general procedure described in Algorithm 2 and the specification for corresponding category in Section 3.3.
Table 3: Cross-domain results on META-DATASET (accuracy (%)).
@@ 4.1 Few-Shot Image Classification Performance
4.2 Cross-Domain Few-Shot Image Classification Performance
To show that ConML is problem-agnostic, we provide learner-agnostic improvement on large-scale cross-domain few-shot image classification problem, obtained on META-DATASET (triantafillou2020meta, ). Table 3 shows the results. The backbone and setting of P>M>F (hu2022pushing, ) is different with the other baselines (triantafillou2020meta, ), so they are not comparable across baselines. ConML is introduced with the same setting as Section 4.1 (inner-task sampling $K=1$ and $\pi_{\kappa}(\mathcal{D}^{\text{tr}}_{\tau}\cup\mathcal{D}^{\text{val}}_{\tau})=\mathcal{D}^{\text{tr}}_{\tau}$, $\phi(a,b)=1-\nicefrac{{a\cdot b}}{{\|a\|\|b\|}}$ (cosine distance) and $\lambda=0.1$.). Note that for P>M>F, ConML is integrated into the meta-training phase, and all other phases remain unchanged. As shown, ConML brings consistent improvement.